40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Sep 7 12:34:54 2017, awade, DailyProgress, Schedule, Task list Sep 7 2017 
    Reply  Mon Sep 11 12:30:19 2017, awade, DailyProgress, Schedule, Task list Sep 11 2017 
Message ID: 1909     Entry time: Mon Sep 11 12:30:19 2017     In reply to: 1899
Author: awade 
Type: DailyProgress 
Category: Schedule 
Subject: Task list Sep 11 2017 

Task list for this week:

Copy paste these check/cross marks for completed/dropped items: ___ ✓ ___ /___ ✗ ___

Kevin/Kira:

  • Calibrate epics channels to Celsius (just using math)
  • Create medm screen, db file and implement PID
  • Design a front and rear panel for the 1U box with the Front Panel Express software (use LIGO-D1700425 and LIGO-D1700426 as templates)
  • Look into good resistive sensor designs for the in vacuum shields controls, talk to Aaron, he's also doing something like this. We have PT100 sensors and REF200 chips, but maybe also look at wheatstone bridges and sensors with higher resistivity.  Make liso files and look at making a noise budget

The PID will be very slow to tune, you may want to ssh into the lab and check periodically.  Also, Yinzi wrote an auto tuning script. You might want to look at that and test it on a smaller thermal mass.

CraIG:

  • FSS LISO PLOTS! TODAY + pdf schematics with "stages" labeled
  • component check and transfer functions of powered up FSS box, we want to be switching out caps and resistors by the end of the week
  • debug of the PSL Frequency Reference Card (21.5 MHz RF for mode cleaners) power regulator issues so we can get this packaged up and working. Make post summarizing failiers so we can keep track of what is broken 
  • check RF power of the 21.5 MHz Ref Card and work out if we need to amplify (because its been splittered), if true then we should package card, mini circuits amplifier, splitters etc into contained 1U box with controls in.  The card has phase control and binary switches on board, so you might work out if you need to come up with a Sub D 9 connector scheme for controlling it.
  • (also awade) gut DC out of back of rack and replace with standard LIGO connector scheme, we can do this while moving the rack

awade:

  • gather quotes on thermal epoxies + specs; masterbond, arctic silver, circuitworks, Lakeshore 7031. Veto non-vacuum compatible ones, want to buy this week
  • Clip for thermal sensors also test soldering methods for kapton heaters, see if flux improves contact, if not examine chemical roughing techniques
  • help craig with FSS box probing and modding
  • package up PMC servo cards in 1u boxes and debug the binary engauge issues that aidan was having
  • ✗ work out how to configure XT1111 16 BIO acromag unit to free up other cards currently being used to drive binary channels
  • debug why startAcromag is not launching as a service, environement variable in .conf script or make a bash wrapper
  • ✓ Order 1u box panels for TTFSS channel controls (remember to add a 5 V indicator to the panel)
  • ✓ Order PCBs for Acromag crate BNC breakouts
  • poke fb1 atf frame builder again and see if living
  • Update LP filters on DAC -> laser input from first order to second order

We also have to find out what needs doing for the pipe repair.

 

ELOG V3.1.3-